OpenStudy (anonymous):

Solve: 2 (5x - 3) - 6x = 2 a. x = -1/4 b. x = 1 c. x = 5/4 d. x = 2

OpenStudy (aroub):

x=2

OpenStudy (lalaly):

10x-6-6x=2 4x=8 x=2

OpenStudy (lalaly):

d

OpenStudy (anonymous):

Okay, so you start like this: 2(5x-3) = 10x - 6 by distribution. 10x - 6 - 6x = 4x - 6 by like terms. 4x - 6 = 2 becomes 4x = 8 by basic algebra. x = 2 by algebra. If this helps, please choose good answer. For free math videos, visit http://www.tutorsean.net